Write the reaction rate expression for the followi
Answer #1

1.

4NH3 + 5O2 => 4 NO+ 6 H2O

rate in terms of disappearance of reactant= - (1/4) d[NH3]/dt = - (1/5)d[O2]/dt

rate in terms of appearance of product= (1/4)d[NO]/dt = (1/6) d[H2O]/dt

2.

as the rate constant is second order, the unit is b) M-1 s-1